Recent Approaches to the Theory of Truth

Andrzej Krzysztof Rogalski

Tarski’s famous paper on the concept of truth (1933) has become a paradigm of all 20^{th} century’s researches on truth. It has greatly influenced the logical semantics, various formal (standard and nonstandard) approaches to linguistics and also the modern logical philosophy. The present paper is focused on:

1) Tarski’s method to interpret his achievements from current perspectives (cf. Peregrin 1999, Rogalski, Bryniarski, Wybraniec-Skardowska 2000),

2) Pointing out shortcomings of the Tarski’s definition of truth and the different new ways of approaches, e.g. Davidson’s approach to meaning by truth-conditions (1984) and recent deflationist formulations by Soames (1999) and, on the other hand, on

3) A non-standard approach, based on the notion of denotation, that can be applied to all languages whose expressions have established meanings. This approach enables to formulate:

a) the syntactic definition of a true sentence which corresponds to the Tarski’s convention (T), and b) the semantic definition of truthfulness of sentences in which the notion of satisfaction in Tarski’s

sense is formally replaced by the use of the notion of denotation (cf. Wybraniec-Skardowska 1998, Rogalski and Wybraniec-Skardowska 2000).

Davidson D. (1984), Inquiries into Truth and Interpretation, Clarendon Press, Oxford. Peregrin J. (1999), Truth and its Nature (if any), Kluwer, Synthese Library, vol. 284. Rogalski A. K. and Wybraniec-Skardowska U. (2000), “On Truth and Denotation”, Proceedings of the Logic Colloquium 1999 (submitted). Rogalski A. K., Bryniarski E., Wybraniec-Skardowska U. (2000), “The Notion of Truth in Systems of Knowledge”, in: Studies in Grammar, Logic and Rethorics (to appear). Soames S. (1999), Understanding Truth, Oxford University Press, New York-Oxford. Tarski A. (1933), Pojecieprawdy w jezykach nauk dedukcyjnych (English translation: The Concept of Truth in Formalized Languages, in: Logic, Semantics, Metamathematics. Papers from 1923 to 1938, Clarendon Press, Oxford 1956). Wybraniec-Skardowska U. (1998), “Logical and Philosophical Ideas in Certain Formal Approaches to Language”, Synthese 116, 231-277.

Address: Faculty of Theology, Lugano, Switzerland

E-mail: andrzej rogalski@hotmail.com